OpenStudy (anonymous):

I know this is very simple but I want to make sure I did it right. The derivative of f(x)= 1/(x^2) is 1 or 1/2?

OpenStudy (datanewb):

The equation: \[f(x) = \frac{1}{x^2} \]can also be written\[f(x) = x^{-2}\]. Taking the derivative of this (the exponent will be -2-1 = -3, and then we multiply by the original exponent to get: \[f'(x) = -2x^{-3}\] Showing this again as a fraction would look like\[f'(x) = \frac{-2}{x^3}\] Note you could have also used the division rule \[\left(\frac{u}{v}\right)' = \frac{u'v-v'u}{v^2} \] This would look like \[f'(x) =\frac{0*x^2-1*2x}{x^4}\\=\frac{-2x}{x^4}\\=\frac{-2}{x^3}\]

OpenStudy (anonymous):

I see, so always flip the fraction when the exponent is in the denominator to work out the derivative?

OpenStudy (datanewb):

Well, you don't have to flip them (for instance you can use the quotient rule for derivatives instead), but I often find it easier to rewrite the equation with the variable in the exponent. Then it is really a simple straight forward process. Thanks for the medal.